OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. Write your equation in point-slope form. Y – 10 = 3/2 (x – 4) ; (4, 2)

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

please note that your slope m, has to satisfy the subsequent condition: \[m*\frac{ 3 }{ 2 }=-1\] so, what is m?

OpenStudy (michele_laino):

please try!

OpenStudy (anonymous):

The slope (m) is 3/2 !

OpenStudy (michele_laino):

sorry, you have to solve the eqaution above in order to find m, please try!

OpenStudy (anonymous):

The slope is 3/2. It's in the equation.

OpenStudy (michele_laino):

I know, but your equation has a different slope, because your line has to be perpendicular not parallel

OpenStudy (anonymous):

The answer is Y – 2 = -2/3 (X – 4) but I don't know how.

OpenStudy (michele_laino):

please solve this equation for m: \[m*\frac{ 3 }{ 2 }=-1\] so you will find your slope m

OpenStudy (anonymous):

I just stated above that I found the answer: Y – 2 = -2/3 (X – 4) all I need to know is how is that the answer?

OpenStudy (michele_laino):

from the equation above we have: \[m=-\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

@Michele_Laino Yes

OpenStudy (michele_laino):

now, please apply this equation: \[y-y _{1}=m(x-x _{1})\] where m is your slope

OpenStudy (anonymous):

Y – 2 = -2/3 (X – 4)

OpenStudy (michele_laino):

and (x_1,y_1) is your point

OpenStudy (michele_laino):

did you get the right answer?

OpenStudy (anonymous):

Yes. Y – 2 = -2/3 (X – 4) is the right answer. What I've been trying to figure out is HOW that is the correct answer?

OpenStudy (michele_laino):

have you understand how?

OpenStudy (michele_laino):

did you understand how to figure out?